OpenStudy (anonymous):

How many solutions does the system have? x = -4y + 4 2x + 8y = 8

OpenStudy (anonymous):

row reduce to echelon form and find out.

OpenStudy (anonymous):

It can have one solution, infinite number of solutions or a unique solution,

OpenStudy (diyadiya):

x = -4y + 4 2x + 8y = 8 first equation can be re-written as x+4y=4 inorder to solve it we multiply this equation by 2 2(x+4y)=2(4) 2x+8y=8 since both equations are same you have infinitely many solutions

OpenStudy (anonymous):

i had actually just worked this out when you both replied. thanks to both of you!

OpenStudy (anonymous):

@Diyadiya I would avoid using the term the "same" in solving system of equations. A better term would be the first equation is a scalar multiple of the second one.

OpenStudy (diyadiya):

@Romero Oh Sorry! Thanks for the correction :)